Two equal charges are located 1m from each other. The force acting between them is 2N. How many Coulombs is each charge?


1
Expert's answer
2022-02-16T08:35:03-0500

Given:

"q_1=q_2=q"

"r=1\:\rm m"

"F=2\:\rm N"


The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}=k\frac{q^2}{r^2}"

Hence, the charge of each point

"q=\sqrt{Fr^2/k}=\sqrt{2*1^2/9*10^9}=1.5*10^{-5}\:\rm C"